OpenStudy (anonymous):

Which of the following is true about the absolute value of -4?

OpenStudy (anonymous):

Which of the following is true about the absolute value of -4?

Parth (parthkohli):

It has distance from 0 as 4, so it is equal to 4.

Parth (parthkohli):

|-4| = 4

OpenStudy (anonymous):

it equals 4

OpenStudy (anonymous):

Which of the following is true about the absolute value of -4? (Points : 2) |-4| = -4 |-4| < -4 |-4| > -4 |-4| = 0

Parth (parthkohli):

Well it is 4 and 4 is greater than -4, so |-4| > -4

OpenStudy (anonymous):

Find the difference of -23 and (32 - 9). 0 18 64 -46

Parth (parthkohli):

-23 and 23 -23 - 23 -46

OpenStudy (anonymous):

Which of the following is equal to |-4 - 1|? 1 - 4 4 - 1 4 + 1 4

OpenStudy (anonymous):

Which of the following is equal to - |-2 +1|? -(-2 + 1) -2 - 1 -2+ 1 -2

Parth (parthkohli):

Why don't you attempt some on your own?

OpenStudy (anonymous):

Trying, but didnt know I had a time limit.

OpenStudy (anonymous):

Had to take medication tonight and can't concentrate.

OpenStudy (anonymous):

Don't usually have this problem

Parth (parthkohli):

|-4 - 1| = |-5| = 5 -|-2 + 1| = -|-1| = -1

OpenStudy (anonymous):

2 min left

Parth (parthkohli):

Just compare the statements.

Parth (parthkohli):

2 min left for?

OpenStudy (anonymous):

Which of the following is equal to |-4 - 1|?

OpenStudy (anonymous):

worksheet

Parth (parthkohli):

Just see...I answered your query right above

OpenStudy (anonymous):

Which of the following is equal to |-4 - 1|? 1 - 4 4 - 1 4 + 1 4

OpenStudy (anonymous):

Find the sum of -23 and (32 - 9). (Points : 2) 0 18 64 9

Parth (parthkohli):

-23 + 23 = ?

OpenStudy (anonymous):

Time is up but thank you for your help.

Parth (parthkohli):

hmm you should have thought a lot more

OpenStudy (anonymous):

i tried

OpenStudy (anonymous):

I am taking a course for photography and don't understand why i have to take math.

OpenStudy (anonymous):

Had a horse accident and developed short term memory and it is hard for me to remember some things that I don't do every day.